OpenStudy (anonymous):

erin and susan both took their dogs for a walk. Together theyw alked 15 miles. If erin walked 3 miles less than susan how far did susan walk with her dog?

OpenStudy (anonymous):

e+s=15 e+3=s

OpenStudy (saifoo.khan):

x+(x-3) = 15

OpenStudy (anonymous):

{e = 6, s = 9}

OpenStudy (anonymous):

is it 10.5?